OpenStudy (anonymous):

Multiply, Give your answer in standard form. (2n2-7n+2)(3n+1) A) 6n3-23n2-n+2 B) 6n3-19n2-n+2 C) 6n3-19n2-13n+2 D)6n3+23n2+13n+2

OpenStudy (anonymous):

=−9n2+3n+2

OpenStudy (anonymous):

So..........it should be D

OpenStudy (anonymous):

Thanks Ill fan you!

OpenStudy (amoodarya):

(2n2-7n+2)(3n+1)= (2n2)(3n+1) + (-7n)(3n+1) +(2)(3n+1) = 6n^3 +2n^2 -21n^2 -7n +6n +2

OpenStudy (anonymous):

Can you also answer this one please... The area o a rectangular painting is given by the trinomial x2+4x-21. What are the possible dimensions of the painting? Use factoring. A) (x+7) and (x+3) B) (x-7) and (x+3) C) (x-7) and (x-3) D) (x+7) and (x-3)

OpenStudy (amoodarya):

x2+4x-21= (x+2)^2-4 -21 =(x+2)^2-25 =(x+2-5) (x+2+5)

OpenStudy (anonymous):

? lol